OpenStudy (anonymous):

19. What is the value of c so that x² – 11x + c is a perfect-square trinomial? 121 –

OpenStudy (anonymous):

ignore the 121 lol

OpenStudy (amistre64):

well, a perfect square is of the form (a+b)^2, which expands out to: a^2+2ab+b^2 lets compare that to what you are given x^2 –11x + c looks like a=x, and 2bx = -11x id say b = -11/2 so c = b^2

OpenStudy (anonymous):

i have worked it out and i cannot find the right answer.. the choices are, 121 121 over 4 -11 over 4 121 over 2

OpenStudy (amistre64):

.... id say b = -11/2 so c = b^2 what praytell is (-11/2)^2 ?

OpenStudy (anonymous):

is this completing the square?

OpenStudy (anonymous):

-11/4?

OpenStudy (amistre64):

no, -11/4 you got the /4 part right now what is (-11)^2 ?

OpenStudy (anonymous):

121/4.. im pretty sure that it!

OpenStudy (amistre64):

\[\frac{-11}{2}*\frac{-11}{2}=\frac{(-11)^2}{(2)^2}\]

OpenStudy (amistre64):

yes, that would be it :)

OpenStudy (anonymous):

okay thank you soo much! XD